Please explain what math was done to get 300 and 400. Also explain how they got 8,820 and 180 please. I circled the parts I'm confused about. Thanks!Dec 10 Dec 11 Dec 12 Dec 19 Carrington Company purchased $10,000 of merchandise from Dansit Company FOB shipping point, terms
Answer #1

Since Cost of Goods Sold for 10000 sale was 4000, so for 1000 sale return, cost of goods sold is 400 i.e 4000/10000 x 1000
Moreover it has scrap value of 100, so loss is 400 -100 = 300

If payment is made within 10 days, there is discount of 2% i.e. 9000 x 2% = 180
Cash = 9000 - 180 = 8820 i.e. cash received net of discount